OpenStudy (anonymous):

I am trying to find help in dividing polynomials when there are two variables : example problem (4x^3y^2+8xy^2-12x^2y^3) / 4xy

OpenStudy (anonymous):

(4x^3y^2+8xy^2-12x^2y^3) / 4xy = 4x^3y^2/ 4xy + 8xy^2 / 4xy - 12x^2y^3 / 4xy = x^2y + 2y - 3xy^2

OpenStudy (anonymous):

How did you split up the variables?

OpenStudy (anonymous):

I understand how to do these problems normally but how do i do it with 2 variables?